OpenStudy (littlet):

the volume of a cylinder is 50 cubic feet. what is the volume of a cone that fits exactly inside the cylinder. provide an explanation and proof to revive full credit

OpenStudy (littlet):

@retirEEd

OpenStudy (littlet):

@zelda can you help?

OpenStudy (anonymous):

What's the equation for volume of a cone, and for volume of a cylinder?

OpenStudy (littlet):

is it (πr2) (h) = πr2h? for cylinder?

OpenStudy (janu16):

so insted aof 50 i used 24 so use this cause i want you to do it on own volume of a cylinder = pi*r^2*h volume of cone = 1/3*pi*r^2*h same radius and height means same r and h given volume of cylinder = 24 so volume of cone = 1/3*24 = 8

OpenStudy (littlet):

Then will my answer be 1/3 * 50= 12.5?

OpenStudy (littlet):

sorry no i meant to divide my 3..

OpenStudy (janu16):

50 divided by 3 is 16.6

OpenStudy (littlet):

1/3*50= 16.67

OpenStudy (janu16):

yess

OpenStudy (littlet):

Thank you!

OpenStudy (janu16):

np

OpenStudy (littlet):

and what would my proof be? @Janu16

OpenStudy (littlet):

what would my proof be?

OpenStudy (littlet):

@mathmale

OpenStudy (janu16):

just explain the steps i wrote

OpenStudy (littlet):

alright thank you again.